OpenStudy (anonymous):

A city’s population, P (in thousands), can be modeled by the equation P=130(1.03)^x , where x is the number of years after January 1, 2000. For what value of x does the model predict that the population of the city will be approximately 170,000? A)8 B)9 C)10 D)11

OpenStudy (anonymous):

P in thousands= 130(1.03)^x 170/130 = 1.03^x \[\ln(\frac{170}{130})=xln(1.03)\] \[\frac{\ln(\frac{170}{130})}{\ln(1.03)}=x\] \[x \approx9\] so the answer is B